OpenStudy (gottennis121):

I'm really brain-dead right now. Please help. Medal! Suzanna wants to put a square fence around her garden. If the garden covers an area of 100 squared, how many feet of fencing does she need? A. 20 B. 40 C. 10 D. 100 I feel like I know the answer, but I don't. Ugg, ever had that feeling before? :/ @e.mccormick @EclipsedStar @ganeshie8 @Hero @TheSmartOne

TheSmartOne (thesmartone):

Do you know what the formula for area is?

OpenStudy (gottennis121):

D is supposed to be 400. Sorry, I'm am half asleep. :/ Formula: L x W

OpenStudy (gottennis121):

So divide? 100/ ?

TheSmartOne (thesmartone):

Exactly! And it says its a square so what would be the formula would be \(\sf Area=side^2\) because width= length

TheSmartOne (thesmartone):

The answer is not 10..

TheSmartOne (thesmartone):

So first we have to find the length of a side. They have told us that the area is 100 so \(\sf area=side^2\) \(\sf 100=side^2\) What is side =?

OpenStudy (gottennis121):

10 then there are 4 sides, so 40?

TheSmartOne (thesmartone):

Correct! So 4x10 gets us the perimeter and so 40 is the correct answer choice! :D

OpenStudy (gottennis121):

Thanks @TheSmartOne

TheSmartOne (thesmartone):

Np :)